OpenStudy (anonymous):

Input the equation of the given line. The line including (4, 7) and (0, 7).

OpenStudy (lieutenantgeneral):

x - y = -7 <-- This equation should be converted into slope-intercept form (y = mx + b). Once you get an equation into y = mx+b form, m = slope, b = y-intercept. So let's convert the equation into y=mx+b form: -y = -7 - x <-- Moved the x to the right side of the equation y = 7 + x <-- Multiplied both sides by -1) y = x + 7 <-- Equation is now in slope-intercept form (y=mx+b). So the slope of the line is 1 because x = 1x. Y-intercept is 7. This means the line intersects the y-axis at coordinate: (0,7). Now to form a line once the y-intercept is plotted on the graph you can: Option 1: Find the x-intercept. The x-intercept is where the line intersects the x-axis. this means y = 0. Plug in y = 0 into y=x+7 and it will give you: 0 = x + 7 <= Plugged in 0 for y) 0 - 7 = x + 7 - 7 <= Subtract 7 to both sides -7 = x or x = -7 <=Simplify So the x intercept is -7, which means the x-intercept coordinate is (-7,0). So plot coordinate (-7,0) on the graph. With the x-intercept (-7,0) and the y-intercept (0,7), use a ruler and draw a straight line through the two points. Option 2: Plug in random x-values into the equation y = x + 7 to get the corresponding y-coordinate. Example, when x = 1, y = 8. When x = 2, y = 9. When x = -1, y = 6. With coordinates (1,8),(2,9) and (-1,6) you can plot these points on a graph and draw a straight line through the points to give your correct line.

OpenStudy (anonymous):

thank you so much for your help!